Problem #1


  1. Complete the following table and identify the quantity that maximizes profit.



Quantity TR MR TC MC


0 0 5

1 10 2

2 20 15

3 30 10

4 10 37




Problem #2


  1. Suppose that each firm in a perfectly competitive market has a short-run total cost of TC = 75 + 500Q – 5Q2 + 0.5Q3, where MC = 500 – 10Q + 1.5Q2.


  1. Calculate the output that minimizes the firm’s AVC.
  2. What is the firm’s shutdown price?
1
Expert's answer
2021-01-21T07:58:53-0500

Problem #1

Quantity TR MR TC MC Profit

0 0 5 -5

1 10 10 7 2 3

2 20 10 15 8 5

3 30 10 25 10 5

4 40 10 37 12 3


The quantities that maximizes profit are 2 and 3 with a maximum profit of 5



Problem #2

"AVC=(500Q-5Q^2+0.5Q^3)\/Q"

"=500-5Q+0.05Q^2"

"AVC=MC"

"500-5Q+0.5Q^2=500-10Q+1.5Q^2"

"-5Q+10Q+0.5Q^2-1.5Q^2=0"

"5Q-Q^2=0"

"Q(5-Q)=0"

"Q=0" and "Q=5"

Q = 5 minimizes the firms cost.

2.shut down price

find TC

"TC=75-50+Q \u2013 5Q^2 + 0.5Q^3"

determine he shut down price by finding the minimum of the AVC

"=( 500Q \u2013 5Q^2 + 0.5Q^3)\/Q"

take the derivative of AVC and set them equal to zero and calculate

"(500-10Q+1.5Q^2)\/Q-(500Q-5Q^2+1.5Q^3)\/Q^2"

"Q=5"

evaluate AVC at Q=5

"500-5(5)+0.05(5)"

"=475.25"

shut down price =475.25